Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Nigth Report 2018-10-30

Observers: M. Berton, E. Congiu
Support Team: S. Benetti, A. Fiore, A. Pastorello
TAT: C. Barbarino, M. Fraser, S. Williams

Calibrations:
Standard EFOSC calibrations - biases, dome flat fields and arcs (evening and morning)

Weather:
Start - Thick clouds (closed)
Middle - Thick clouds (closed)
End - Thin clouds, seeing ~1.3

Instrument - EFOSC

UT            Target               Type                Setup                  airmass             Seeing              notes
-----------------------------------------------------------------------------------------
07:00        image analysis
07:23        AT2018hmy      class                gr13                    1.39                    1.3                    Ia
07:45        SN2018gsk       follow              gr11                     1.38                   1.2
08:26        LTT3864            STD                 gr11+gr13+gr16  1.65                   1.6                    
       

               



Comments